OpenStudy (anonymous):

[4_06]Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (anonymous):

The formula for a line with coordinates is: \[y-y_{1}=m(x-x_{1})\] Since they give you the gradient and the coordinates just plug them in to the formula: y - (-1) = 4( x - (-3) ) y + 1 = 4 ( x + 3 )